# Robust Parameter Estimation for Rational Ordinary Differential Equations

### 05 Dec 2023 v2

@@ Abstract
In a typical existing approach, one first tries to make a good initial guess for the parameter values. Then, in a loop, the corresponding outputs are computed by solving the ODE numerically, followed by computing the error from the given time series data. If the error is small, the loop terminates and the parameter values are returned. Otherwise, heuristics/theories are used to possibly improve the guess and continue the loop. A downside of this approach is non-robustness, as there are no guarantees for the result of the loop iterations to be predictably close to the true parameter values.
In this paper, we propose a new approach, which does not suffer from the above non-robustness. In particular, it does not require making good initial guesses for the parameter values. Instead, it uses differential algebra, interpolation of the data using rational functions, and multivariate polynomial system solving, and has a potential for a complete user control over the error of the estimation (the actual error analysis is left for the future research). We also compare the performance of the resulting software with several other estimation software packages.
@@ 1.2 State of the art
Due to its importance, this problem has been the subject of intensive research efforts that have yielded various theories and methods. They can be roughly categorized into three approaches: shooting, two-stage, and algebraic.
Shooting approaches (also known as simulation-based approaches) directly use the ODE solver to compute the loss function of the current estimate. This is used to define a nonlinear optimization to improve the parameter estimates. This popular approach has seen many software implementations; to list a few: AMIGO2, COPASI, Data2Dynamics, SBtoolbox2/IQM; see [3, 5, 38, 65, 58, 56, 36, 29, 23, 21, 2, 54, 25, 63, 19]. Many studies have further developed the shooting methods to improve robustness, adding techniques such as multiple shooting [10, 1, 60], automatic differentiation and adjoint techniques for more accurate derivatives [41, 46, 33], and landscape smoothing techniques like the prediction-error method [34], all of which help the optimization techniques avoid local minima.
Two-stage approaches (also known as collocation, smoothing, or principle differential analysis) avoid numerically solving the ODE by fitting collocation polynomials to the data to obtain predictions of the data which are then directly fit against the data [52, 14]. As the result, these methods are generally much more efficient. Previous studies have used various mathematical objects such as B-splines [13, 48, 35] and local least-squares [37, 55, 53, 15]. One downside to two-stage methods is requiring observation of all states. Furthermore, it has been noted that the error introduced through the derivative estimation tends to lead to less accurate final results in real-world scenarios with less dense data [48, 14]. These studies suggest finishing a parameter estimation by using the output of a two-stage method as the initial condition to a simulation-based method.
@@
1.3 Our approach and novelty
Our approach could roughly be classified as a two stage approach. However, it is substantially different from the previous two stage approaches, in that it computes optimal parameters robustly. For instance, it can reliably handle even degenerate cases such as when the parameters are only locally identifiable. It achieves the robustness by judiciously combining:
@@ 1.3 Our approach and novelty
Our approach has the following novel features:
  1. 1.
    Our approach does not involve making initial guesses or global optimization; it instead relies on differential algebraic properties of the ODE model to propose candidates for parameter values.
  2. 2.
    Out of the possible solution candidates, the procedure computes the best fit for the given data including cases in which such fit may not be unique. That is, accounting for possible local identifiability of the model, we report all valid parameter values. The user can then quickly filter out some (or all) of the not possible parameter values if the users has additional information about these values such as ranges they take values in.
As such, this technique can be used in isolation for dense data, or together with shooting methods by providing more robust initial conditions for sparse data.
@@ 1.4 Software Implementation
We implemented our approach into a Julia package called “ParameterEstimation.jl”. All code and benchmarking data are available at https://github.com/iliailmer/ParameterEstimation.jl.
@@ 2 Problem Statement
In this section, we precisely state the problem solved by our approach and illustrate it using a toy example. We also show how to enter/solve the problem using our soft implementation.
@@
3 Our Approach
We will briefly describe our approach by first giving a flow chart of the steps.
@@ 3 Our Approach
    Toy example: There is only one value for the parameters.
@@ 4 Implementation details
Our software supports a variety of modern interpolation methods, via BaryRational.jl and other interpolation packages. Of note, we support Fourier extension, Floater-Hormann rational interpolation [24], as well as the “AAA" algorithm introduced in [45], which uses a barycentric rational representation adapted to input data. Experimental data show that AAA rational interpolation interpolates smooth functions in such a way that, for a wide variety of data outputs from differential equations, even higher order derivatives are well interpolated. In our program, we use AAA as a default interpolation scheme, but allow iterating over multiple schemes in a parallel fashion to select the best one.
@@
We list a few limitations of the current implementation:
  • As a two-stage method, the issues common to this class of methods apply. The data must be time series data which is sufficiently dense for the rational polynomial fitting to give accurate derivative approximations.
  • Our program supports rational ODEs models (the right-hand sides are rational functions), while software such as AMIGO2 and IQM also support ODE models that are not necessarily rational.
@@ 5.2 Settings used for Software Packages
IQM, SciML and AMIGO2 require search ranges or initial guesses for the parameters and initial conditions or certain additional settings to be specified by the user that pertain specifically to the optimization task performed by each program, for instance, maximal # of iterations for gradient-based optimization, specific loss/error functions to be optimized.
@@ 5.5 Discussion
We make the following observations from the test results.
  • Relative error: For most of the scenarios considered here, our system was able to outperform in estimation error. In all but one instance, we observed that our method outperformed the others with regard to estimation error.
  • Multiple solutions: Our program’s advantage is its ability to recover multiple solutions if appropriate (local identifiability) while optimization-based programs may miss a valid value of a parameter.
One of the most notable improvements of our approach over the shooting-based methods is the lack of required hyperparameters or the necessity of choosing the search range for the parameter values. In contrast, e.g., in AMIGO2, one must choose lower and upper bounds for the ODE parameters and initial conditions. In particular, in the Crauste Model, eq. 14, if the upper bound of the range for all parameters is increased from 1.0 to 2.0, the median (mean) of RMSRE’s increased from 0.0% (2.6%) to 77.6% (105.7%). However, if the range is further increased to 3.0, the median (mean)of RMSRE’s decreased to 24.5% (51.7%). (See Random Data 3, 5, and 8-10 in the Crauste table of Appendix for specific instances of this behavior.) This shows the non-monotonic behavior of tuning the approaches that rely on global optimization, which can add difficulty for scientists to accurately use the methods in practice. This is not a criticism of these approaches but is rather a note that the potential users have to be aware of.
Our parameter estimation approach does not require any hyperparameters while giving all solutions (e.g., if some parameters are only locally but not globally identifiable), making it easier for a practicing scientist to quickly understand the full landscape of solutions.
@@ 6 List of Differential Models
In this section, we present the ODE systems used in our benchmarks.
@@
$\begin{cases}y_{1}=x_{1}\\ y_{2}=x_{2}\end{cases}$ (3)
@@
$\begin{cases}y_{1}=x_{1}\\ y_{2}=x_{2}\end{cases}$ (5)
@@
$\begin{cases}y_{1}=w,~{}y_{2}=z\\ y_{3}=x,~{}y_{4}=y+v\end{cases}$ (9)
@@
$\begin{cases}y_{1}=x_{1}\\ y_{2}=x_{2}\end{cases}$ (11)
@@
$\begin{cases}y_{1}=N,~{}y_{2}=E\\ y_{3}=S+M,y_{4}=P\end{cases}$ (15)
@@
$\begin{cases}y_{1}=x_{4}\\ y_{2}=x_{5}\end{cases}$ (17)
@@
$\begin{cases}y_{1}=x_{1}\\ y_{2}=x_{2}\\ y_{3}=x_{3}+x_{4}\end{cases}$ (19)
@@
$\begin{cases}y_{1}=I\\ y_{2}=N\end{cases}$ (21)
@@ Acknowledgements
The authors are grateful to the CCiS at Queens College and CIMS NYU for the computational resources, to Julio Banga and Eva Balsa-Canto for extensive help and feedback on the use of AMIGO2. The authors are also grateful to Xinglong Zhou and Jianing Qi for initial explorations of the role of polynomial interpolation in this problem. The authors are also grateful to Henning Schmidt, developer of IQRTools/IQDesktop for numerous discussions while we tried to install the software. We thank Gleb Pogudin for helpful comments on our algorithm and additional information on other parameter estimation methods.

### 17 Dec 2023 v3

@@ Abstract
In typical existing approaches, an initial guess for the parameter values is made from a given search interval. Then, in a loop, the corresponding outputs are computed by solving the ODE numerically, followed by computing the error from the given time series data. If the error is small, the loop terminates and the parameter values are returned. Otherwise, heuristics/theories are used to possibly improve the guess and continue the loop.
These approaches tend to be non-robust in the sense that their accuracy depend on the search interval and the true parameter values; furthermore, they cannot handle the case where the parameters are locally identifiable.
In this paper, we propose a new approach, which does not suffer from the above non-robustness. In particular, it does not require making good initial guesses for the parameter values or specifying search intervals. Instead, it uses differential algebra, interpolation of the data using rational functions, and multivariate polynomial system solving. We also compare the performance of the resulting software with several other estimation software packages.
@@ 1.2 State of the art
Due to its importance, this problem has been the subject of intensive research efforts that have yielded various theories and approaches. They can be roughly categorized into three approaches: shooting, two-stage, and algebraic.
Shooting approaches (also known as simulation-based approaches) directly use the ODE solver to compute the loss function of the current estimate. This is used to define a nonlinear optimization to improve the parameter estimates. This popular approach has seen many software implementations; to list a few: AMIGO2, COPASI, Data2Dynamics, SBtoolbox2/IQM; see [3, 5, 38, 65, 58, 56, 36, 29, 23, 21, 2, 54, 25, 63, 19]. Many studies have further developed the shooting approaches to improve robustness, adding techniques such as multiple shooting [10, 1, 60], automatic differentiation and adjoint techniques for more accurate derivatives [41, 46, 33], and landscape smoothing techniques like the prediction-error method [34], all of which help the optimization techniques avoid local minima.
Two-stage approaches (also known as collocation, smoothing, or principle differential analysis) avoid numerically solving the ODE by fitting collocation polynomials to the data to obtain predictions of the data which are then directly fit against the data [52, 14]. As the result, these approaches are generally much more efficient. Previous studies have used various objects such as B-splines [13, 48, 35] and local least-squares [37, 55, 53, 15]. One downside to two-stage approaches is requiring observation of all states. Furthermore, it has been noted that the error introduced through the derivative estimation tends to lead to less accurate final results in real-world scenarios with less dense data [48, 14]. These studies suggest finishing a parameter estimation by using the output of a two-stage approach as the initial condition to a simulation-based approach.
@@
1.3 The proposed approach and novelty
The proposed approach could roughly be classified as a two stage approach. However, it is substantially different from the previous two stage approaches, in that it estimates parameters robustly, in the following sense:
  1. 1.
    The proposed approach produces good estimates no matter what the specified search intervals are. In fact, it does not even require specifying any search interval.
  2. 2.
    The proposed approach produces good estimates no matter where the true parameter values are.
  3. 3.
    The proposed approach produces good estimates even when the parameters are not globally identifiable, as long as they are locally identifiable.
It achieves the robustness by judiciously combining:
@@ 1.3 The proposed approach and novelty
As such, this technique can be used in isolation for dense data, or together with shooting approaches by providing more robust initial conditions for sparse data.
@@ 1.4 Software Implementation
We implemented the proposed approach into a Julia package called “ParameterEstimation.jl”. All code and benchmarking data are available at https://github.com/iliailmer/ParameterEstimation.jl.
@@ 2 Problem Statement
In this section, we precisely state the problem solved by the proposed approach and illustrate it using a toy example. We also show how to enter/solve the problem using our soft implementation.
@@
3 The proposed Approach
We will briefly describe the proposed approach by first giving a flow chart of the steps.
@@ 3 The proposed Approach
    Toy example:
    There is only one value for the parameters.
@@ 4 Implementation details
Our software supports a variety of modern interpolation methods, via BaryRational.jl and other interpolation packages. Of note, we support Fourier extension, Floater-Hormann rational interpolation [24], as well as the “AAA" algorithm introduced in [45], which uses a barycentric rational representation adapted to input data. Experimental data show that AAA rational interpolation interpolates smooth functions in such a way that, for a wide variety of data outputs from differential equations, even higher order derivatives are well interpolated. In our software, we use AAA as a default interpolation scheme, but allow iterating over multiple schemes in a parallel fashion to select the best one.
@@
We list a few limitations of the current preliminary implementation:
  • As a two-stage approach, the issues common to this class of approaches apply. The data must be time series data which is sufficiently dense for the rational polynomial fitting to give accurate derivative approximations.
  • Our software supports rational ODEs models (the right-hand sides are rational functions), while software such as AMIGO2 and IQM also support ODE models that are not necessarily rational.
@@ 5.2 Settings used for Software Packages
IQM, SciML and AMIGO2 require search ranges or initial guesses for the parameters and initial conditions or certain additional settings to be specified by the user that pertain specifically to the optimization task performed by each software package, for instance, maximal # of iterations for gradient-based optimization, specific loss/error functions to be optimized.
@@ 5.5 Discussion
The inspection of the test results in Table 1 and Appendix shows that the proposed approach outperforms the others with regard to estimation error in all but one instance. Furthermore, the proposed approach is robust in the following aspects:
Search Interval: See Table 1. Note that IQM, SciML and AMIGO2 (based on shooting approaches) require the user to specify a search interval. Note also that their relative errors are dependent on search interval, sometimes even non-monotonically. For an example, see the relative error of AIMOG2 on the Crauste Model. If the upper bound of the search interval is increased from 1.0 to 2.0, the median (mean) of the relative error is increased from 0.0% (2.6%) to 77.6% (105.7%). However, if the upper bound is increased to 3.0, the median (mean) of the relative error is decreased to 24.5% (51.7%). See Random Data 3, 5, 9 and 10 in the Crauste table of Appendix for specific instances of this behavior. This non-monotonic behavior could add difficulty to the users in specifying a suitable search interval.
The proposed approach’s relative error is, by design, independent of search interval, because it does not require/use search interval.
Parameter values: See the tables in Appendix. Note that the relative errors of IQM, SciML and AMIGO2 are dependent on the values of the parameters (randomly sampled from [0.1, 0.9]), sometime drastically. For an example, see the Crauste table. AMIGO2 has essentially no error for some random values of parameters but huge error for the others.
The proposed approach’s relative error is practically independent of the values of the parameters.
Multiple solutions: See Table 1. The first seven models are globally identifiable (hence with unique solution for the parameters). The last three models contain parameters that are only locally identifiable (hence with multiple solutions for the parameter values). Note that the relative errors of IQM, SciML and AMIGO2 are significantly larger for locally identifiable models. It is because the optimization-based approaches may miss an intended value of a parameter.
The proposed approach’s relative error is practically independent of identifiability, as long as it is at least locally identifiable. In fact, it finds all the solutions, making it easier for the users to quickly understand the full landscape of solutions.
@@ 6 List of Differential Models
In this section, we present the ODE systems used in the benchmarks.
@@
$\begin{cases}y_{1}=x_{1},\\ y_{2}=x_{2}\end{cases}$ (3)
@@
$\begin{cases}y_{1}=x_{1},\\ y_{2}=x_{2}\end{cases}$ (5)
@@
$\begin{cases}y_{1}=w,\\ y_{2}=z,\\ y_{3}=x,\\ y_{4}=y+v\end{cases}$ (9)
@@
$\begin{cases}y_{1}=x_{1},\\ y_{2}=x_{2}\end{cases}$ (11)
@@
$\begin{cases}y_{1}=N,\\ y_{2}=E,\\ y_{3}=S+M,\\ y_{4}=P\end{cases}$ (15)
@@
$\begin{cases}y_{1}=x_{4},\\ y_{2}=x_{5}\end{cases}$ (17)
@@
$\begin{cases}y_{1}=x_{1},\\ y_{2}=x_{2},\\ y_{3}=x_{3}+x_{4}\end{cases}$ (19)
@@
$\begin{cases}y_{1}=I,\\ y_{2}=N\end{cases}$ (21)
@@ Acknowledgements
We, the authors, are grateful to Julio Banga and Eva Balsa-Canto for extensive feedback on the use of AMIGO2, to Henning Schmidt, developer of IQRTools/IQDesktop for numerous discussions on installation of the software, to Gleb Pogudin for helpful feedback on the proposed algorithm and information on other parameter estimation approaches, to Xinglong Zhou and Jianing Qi for initial explorations of the role of polynomial interpolation, and finally to the CCiS at Queens College and CIMS NYU for the computational resources.